# AutoRE: Document-Level Relation Extraction with Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional abilities in comprehending and generating text, motivating numerous researchers to utilize them for Information Extraction (IE) purposes, including Relation Extraction (RE). Nonetheless, most existing methods are predominantly designed for Sentence-level Relation Extraction (SentRE) tasks, which typically encompass a restricted set of relations and triplet facts within a single sentence. Furthermore, certain approaches resort to treating relations as candidate choices integrated into prompt templates, leading to inefficient processing and suboptimal performance when tackling Document-Level Relation Extraction (DocRE) tasks, which entail handling multiple relations and triplet facts distributed across a given document, posing distinct challenges. To overcome these limitations, we introduce AutoRE, an end-to-end DocRE model that adopts a novel RE extraction paradigm named RHF (Relation-Head-Facts). Unlike existing approaches, AutoRE does not rely on the assumption of known relation options, making it more reflective of real-world scenarios. Additionally, we have developed an easily extensible RE framework using a Parameters Efficient Fine Tuning (PEFT) algorithm (QLoRA). Our experiments on the RE-DocRED dataset showcase AutoRE’s best performance, achieving state-of-the-art results, surpassing TAG by 10.03% and 9.03% respectively on the dev and test set. The code is available [^1] https://github.com/bigdante/AutoRE and the demonstration video is provided [^2] https://www.youtube.com/watch?v=IhKRsZUAxKk.
footnotetext: ${}^{\dagger}$ equal contribution
@@ 1 Introduction
The rise of LLMs, such as GPT-4 (Achiam et al., 2023) and Llama2 (Touvron et al., 2023), has significantly propelled the progress of Natural Language Processing due to their strong capabilities in text understanding, generation, and generalization (Zhao et al., 2023). There has been an increasing interest in using LLMs to generate structured information for IE tasks (Xu et al., 2023; Li et al., 2023b; Wadhwa et al., 2023), and making impressive progress. Typical IE tasks using LLMs include Named Entity Recognition (NER) (Wang et al., 2023a), Relation Extraction (RE) (Zhou et al., 2023), and Event Extraction (EE) (Xu et al., 2023; Huang et al., 2023). Despite the outstanding result, the performance of current LLMs in RE is still far from satisfactory.
@@
Figure 1: The result on Re-DocRED. AutoRE(-A), for different PLMs, achieves SoTA.
Underperformance in DocRE Tasks. We evaluated several high-performing LLMs on the DocRE task, specifically using the test set of Re-DocRED (Tan et al., 2022). The models included GPT-3.5-turbo  [^3] https://chat.openai.com/chat(ChatGPT), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) (Mistral-7B), Vicuna-7B-v1.5 (Chiang et al., 2023) (Vicuna-7B), and ChatGLM3-6B (Du et al., 2022). Our results indicate that, without specific fine-tuning, the performance of these language models on DocRE tasks is suboptimal, as shown in the blue bars in Figure 1.
@@
We have innovated a new pipeline RE paradigm called RHF to address challenges identified in existing RE paradigms. We comprehensively redefined the 96 relation descriptions and crafted simplified relation extraction templates, developing an instructive tuning dataset based on Re-DocRED. Utilizing the Mistral-7B model and applying PEFT (Parameter Efficient Fine-Tuning) using QLoRA Dettmers et al. (2023), our model achieved state-of-the-art (SOTA) performance on the Re-DocRED test set. Key contributions include:
@@ 2 Related work
LLMs for DocRE. Researchers have been employing LLMs to tackle RE tasks. For example, ChatIE (Wei et al., 2023) deconstructs the complex RE process into assembling the outputs from multiple rounds of Question-Answer into a final structured format. However, the performance of LLMs in RE tasks still lags behind SOTA models. Han et al. concluded that ChatGPT does not adequately comprehend the subject-object relationships in RE tasks. Similarly, Li et al. noted that in Standard-IE settings, ChatGPT’s performance is generally not as effective as BERT-based models. Moreover, most models are tested on SentRE. To test LLMs for DocRE, we conducted tests on ChatGPT, Mistral-7B, Vicuna-7B, and ChatGLM3-6B and revealed that the current performance is far from satisfactory, as illustrated in Figure 1. This aligns with findings reported by (Li et al., 2023c), indicating that current models still fall significantly short in performance on DocRE.
RE Prompt Template. The models fine-tuned on LLMs for RE operate on a prompt-based or instruction-driven mechanism(Beurer-Kellner et al., 2023), engaging in a question-and-answer format to execute RE tasks. ChatIE (Wei et al., 2023), InstructUIE (Wang et al., 2023b), and YAYI Xiao et al. (2023) while demonstrating formidable capabilities in IE, exhibit considerable limitations in their RE prompt templates. A common method in their RE process involves embedding a list of relations into the model’s prompt template as alternatives. However, this approach becomes impractical when dealing with DocRE, such as the 96 relations in the Re-DocRED. This limitation is acknowledged by Wadhwa et al., who concludes that “for datasets with long texts or a large number of targets, it is not possible to fit detailed instructions in the prompt”.
@@ 3.1 RE Paradigms
Figure 2: Processing step of different RE paradigms.
We test these paradigms with ChatGPT and the results are displayed in Table 1, and the specific prompt is included in the Appendix A.1. We arrived at the following conclusion:
@@ 3.2 Dataset Processing
We used the Re-DocRED dataset for fine-tuning, refining it by removing duplicates and ensuring factual accuracy. This involved adjusting reciprocal relations like ’’follows’’ and ’’followed by’’ to accurately represent inversion, enhancing the dataset’s robustness and precision.
@@
Addressing this, we systematically rewrote all 96 relation descriptions, markedly improving model performance (Table 2). An example of our revised description is as follows. (Details of all relation descriptions are presented in Appendix A.2).
@@
Finally, in line with the RHF paradigm, we crafted instruction fine-tuning templates, breaking down the RE process of each sample into three distinct steps. The specific details of these templates can be found in the Appendix under A.3.
@@ 4.1 Experimental Setup
4.2 Overall Result
After training three distinct QLoRA modules, we test the performance on the Re-DocRED and then combine three QLoRAs to get the final performance on the dev set and test set, the result is shown in Table 3. When compared with TAG (Zhang et al., 2023) as a baseline which firstly reported the end-to-end RE on Re-DocRED, our method has achieved SoTA results, as shown in Table 4. In both the dev set and test set, the performance improvement of AutoRE over TAG is approximately 7.44% on the dev set, and about 5.12% on the test set demonstrating the effectiveness of our approach. Furthermore, by decomposing the task into three subtasks and training with three LoRA modules, we not only achieved excellent results but also naturally acquired an easily extendable trait. This allows for targeted improvement of a specific module’s performance without impacting the performance of other subtasks. Additionally, it is worth noting that our work is the first to utilize large language models for processing the Re-DocRED dataset. AutoRE can serve as a reference for subsequent research in this field.
@@ 4.3 Ablation
In the ablation study, we employed Mistral-7B to fine-tune the paradigms mentioned before, revealing that the RHF model yields the best performance when solely utilizing one QLoRA module. This finding substantiates our initial hypothesis during paradigm selection: employing a step-by-step approach enhances the extraction of triplet facts while significantly reducing erroneous triplets through fine-tuning. Building on this, we compared the impact of including descriptions versus omitting them. The results confirmed that incorporating proper relation descriptions indeed benefits the model, as shown in 4. Additionally, we explored the effectiveness of training the entire dataset with one QLoRA versus independently training different stages of RHF with three distinct QLoRAs. The latter approach demonstrated superior performance. We believe this is due to the data imbalance among predicting relations, predicting head entities, and predicting factual triples in the RHF paradigm, with the data volume for the three subtasks begin approximately 2.8%, 24.23%, and 72.97%, respectively. When combined, the model tends to favor the prediction of triples, while its capability to predict relations is relatively insufficient.
Additionally, we have applied this framework to Vicuna-7B and ChatGLM3-6B, and both models surpassed the current SOTA levels, demonstrating the universality of AutoRE framework. The comparative results of these experiments are illustrated in the accompanying Figure 4. Vicuna-7B scored the highest, whereas ChatGLM3-6B was somewhat lower. This might be due to ChatGLM3-6B having a higher proportion of Chinese in its pre-training, while it was tested on an English task. Now, we have deployed the system on the online platform  [^5] https://models.aminer.cn/neptune/ for users to access and experience, as shown in Figure 3.
@@ 5 Conclusion
Appendix A Appendix
A.1 ChatGPT prompt
We provide testing prompts for the Re-DocRED dataset under different paradigms using ChatGPT in Table 5. For brevity, we only provide two representative relation extraction prompt words. The rest are similar to these.
A.2 Relation Description
We display the new refined relation descriptions provided for each relation in Table 6.
A.3 Instruct Template
In Table 7, we provide a display of how we constructed our training data using simple prompt templates, the extraction of relations, the extraction of the head entity, and finally the triplet extraction.
Model dev F1 test F1
TAG 49.34 49.38
AutoRE-ChatGLM3-6B 49.86 51.11
AutoRE-Mistral-7B 53.01 51.91
AutoRE-Vicuna-7B 54.29 53.84
Table 4: The results of AutoREs for different PLMs. Compared with TAG, AutoREs all achieve SoTA.
[…]
[width=.85]figures/Ablation.pdf
Paradigms Prompt
D-R-F
Given a passage: {sentences}, and relation list: {relation_list}
Check the passage, and find which relations can be derived from the passage.
Your output format is as following:
relation1
relation2
one example like:
country of citizenship
father
The relations must be in the relation list.
If no relation in the sentence, you should only output:
no relation
Given a relation: {relation}.
Provided a passage: "{sentences}".
Derive all the triplet facts from the passage according to the given relations.
Your output format is as following:
["subject",{relation},"object"]
["subject",{relation},"object"]
The subject and object should be entity from the passage.
D-R-H-F
Given a passage: {sentences}, and relation list: {relation_list}
Check the passage, and find which relations can be derived from the passage.
Your output format is as following:
relation1
relation2
one example like:
country of citizenship
father
The relations must be in the relation list.
If no relation in the sentence, you should only output:
no relation
Given the relation: {relation}.
Now the passage is: {sentences}.
Derive all the entity from the passage that can serve as the subject of the {relation}.
Your output format is as following:
entity1
entity2
The entities should all be from the passage.
Given the relation: {relation}.
Now the passage is: {sentences}.
Derive all the triplet facts from the passage that take {subject} as subject.
Your output format is as following:
[{subject},{relation},object]
[{subject},{relation},object]
The object should be entity from the passage.
Table 5: ChatGPT prompt template for RE on Re-DocRED.
Relation Description
located in the administrative territorial entity
In the ’located in the administrative territorial entity’ relation, the subject, a place, event, or item, resides or takes place in the object, an administrative region. Example: (Harvard University, located in the administrative territorial entity, Cambridge, Massachusetts).
country
For the ’country’ relation, the subject pertains to a non-human entity, such as an organization, place, or event. The object signifies the sovereign state where the subject is based or occurs. Example: (Amazon Inc, country, United States).
country of citizenship
The ’country of citizenship’ relation denotes that the subject, an individual, is recognized as a citizen by the object, a country. Example: (Elon Musk, country of citizenship, United States).
contains administrative territorial entity
The relation ’contains administrative territorial entity’ involves a subject, an administrative territory, encompassing the object, a subdivision or part of this administrative territory. Example: (California, contains administrative territorial entity, Los Angeles).
has part
The ’has part’ relation reflects that the subject, an entity or whole, comprises the object, a part or component of the subject. Example: (A car, has part, engine).
date of birth
In the ’date of birth’ relation, the subject, a person, was born on the object, the specified date. Example: (John Doe, date of birth, January 1, 1990).
part of
In the ’part of’ relation, the subject, a component or section, belongs to the object, a larger whole or aggregate. Example: (Engine, part of, a car).
notable work
The ’notable work’ relation indicates a significant work assigned to the subject, a creator, while the object is that noted scientific, artistic, or literary work itself. Example: (Jane Austen, notable work, Pride and Prejudice).
publication date
The ’publication date’ relation marks when the subject, a work, was first published or released, with the object being that specific date. Example: (Pride and Prejudice, publication date, 1813).
inception
In the ’inception’ relation, the subject, an event or a item (not a person), came into existence at the object, a specific date or point in time. Example: (Google, inception, September 4, 1998).
date of death
The ’date of death’ relation specifies when the subject, a once-living person, died. The object is the particular date of demise. Example: (Albert Einstein, date of death, April 18, 1955).
Table 6: New designed relation descriptions. We only present part of the descriptions of 96 relations. The whole relation descriptions can be found via this link: https://github.com/bigdante/AutoRE.
Submission Instruct Tuning Template
relation_template
Given a passage: "{sentences}" List any underlying relations..
entity_template
Given a relation "{relation}", and its description: "{description}" and a passage: "{sentences}", list entities that can be identified as suitable subjects for the relation.
fact_template
"Given relation "{relation}" and relation description: "{description}".Provided a passage: "{sentences}"List all triple facts that take "{relation}" as the relation and "{subject}" as the subject.
Table 7: Instruct Tuning Template for RHF.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional abilities in comprehending and generating text, motivating numerous researchers to utilize them for Information Extraction (IE) purposes, including Relation Extraction (RE). Nonetheless, most existing methods are predominantly designed for Sentence-level Relation Extraction (SentRE) tasks, which typically encompass a restricted set of relations and triplet facts within a single sentence. Furthermore, certain approaches resort to treating relations as candidate choices integrated into prompt templates, leading to inefficient processing and suboptimal performance when tackling Document-Level Relation Extraction (DocRE) tasks, which entail handling multiple relations and triplet facts distributed across a given document, posing distinct challenges. To overcome these limitations, we introduce AutoRE, an end-to-end DocRE model that adopts a novel RE extraction paradigm named RHF (Relation-Head-Facts). Unlike existing approaches, AutoRE does not rely on the assumption of known relation options, making it more reflective of real-world scenarios. Additionally, we have developed an easily extensible RE framework using a Parameters Efficient Fine Tuning (PEFT) algorithm (QLoRA). Our experiments on the RE-DocRED dataset showcase AutoRE’s best performance, achieving state-of-the-art results, surpassing TAG by 10.03% and 9.03% respectively on the dev and test set. The code is available [^1] https://github.com/THUDM/AutoRE and the demonstration video is provided [^2] https://www.youtube.com/watch?v=IhKRsZUAxKk.
footnotetext: LX and DZ contributed equally.footnotetext: JT is the corresponding author.
@@ 1 Introduction
The rise of LLMs, such as GPT-4 (Achiam et al., 2023) and Llama2 (Touvron et al., 2023), has significantly propelled the progress of natural language processing due to their strong capabilities in text understanding, generation, scientific reasoning (Zhang et al., 2024a, b), social bot detection (Zhou et al., 2024), and generalization (Zhao et al., 2023). There has been an increasing interest in using LLMs to generate structured information for IE tasks (Xu et al., 2023; Wadhwa et al., 2023) and making impressive progress. Typical IE tasks using LLMs include Named Entity Recognition (NER) (Wang et al., 2023a), Relation Extraction (RE) (Zhou et al., 2023), and Event Extraction (EE) (Xu et al., 2023). Despite the outstanding result, the performance of current LLMs in RE is still far from satisfactory.
@@
Figure 1: The result on the test set of Re-DocRED. AutoRE (-A) achieves SOTA for different LLMs.
Underperformance on DocRE Tasks. We evaluated several high-performing LLMs on the document-level RE (DocRE) task, specifically using the test set of Re-DocRED (Tan et al., 2022). These models included GPT-3.5-turbo [^3] https://chat.openai.com/chat(ChatGPT), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) (Mistral-7B), Vicuna-7B-v1.5 (Chiang et al., 2023) (Vicuna-7B), and ChatGLM3-6B (Du et al., 2022). Our results indicate that, without specific fine-tuning, the performance of these language models on DocRE tasks is suboptimal, as shown in the blue bars in Figure 1.
@@
To address challenges identified in existing RE paradigms, we innovate a new pipeline RE paradigm, Relation-Head-Facts (RHF). We comprehensively redefined the 96 relation descriptions and crafted simplified relation extraction templates, developing an instruction-tuning dataset based on Re-DocRED. Utilizing the Mistral-7B model with Parameter Efficient Fine-Tuning (PEFT), QLoRA Dettmers et al. (2023), our model achieved state-of-the-art (SOTA) performance on the Re-DocRED test set. Key contributions of our work include:
@@ 2 Related work
LLMs for DocRE. Researchers have been employing LLMs to tackle RE tasks. For example, ChatIE (Wei et al., 2023) deconstructs the complex RE process into assembling the outputs from multiple rounds of Question-Answer into a final structured format. PromptRE (Gao et al., 2023) integrates LLM prompting with data programming to deal with DocRE. However, the performance of LLMs on RE tasks still lags behind SOTA models. Han et al. concluded that ChatGPT does not adequately comprehend the subject-object relationships in RE tasks. Similarly, Li et al. noted that in Standard-IE settings, ChatGPT’s performance is generally not as effective as BERT-based models. Moreover, most models are tested on SentRE. To test LLMs for DocRE, we conducted tests on ChatGPT, Mistral-7B, Vicuna-7B, and ChatGLM3-6B and revealed that the current performance is far from satisfactory, as illustrated in Figure 1. This aligns with findings reported by (Li et al., 2023b), indicating that current models still fall significantly short in performance on DocRE.
RE Prompt Template. These models fine-tuned on LLMs for RE operate on a prompt-based or instruction-driven mechanism (Beurer-Kellner et al., 2023), engaging in a question-and-answer format to execute RE tasks. ChatIE (Wei et al., 2023), InstructUIE (Wang et al., 2023b), and YAYI Xiao et al. (2023) while demonstrating formidable capabilities in IE, exhibit considerable limitations in their RE prompt templates. A common method in their RE process involves embedding a list of relations into the model’s prompt template as alternatives. However, this approach becomes impractical when dealing with DocRE, such as the 96 relations in the Re-DocRED. This limitation is acknowledged by Wadhwa et al., who concludes that “for datasets with long texts or a large number of targets, it is not possible to fit detailed instructions in the prompt”.
@@ 3.1 RE Paradigms
Figure 2: Processing steps of different RE paradigms.
We test these paradigms with ChatGPT and the results are displayed in Table 1. We provide testing prompts for the Re-DocRED dataset under different paradigms using ChatGPT in Table 5. For brevity, we only provide two representative relation extraction prompt words. The rest are similar to these. We arrived at the following conclusion:
@@ 3.2 Dataset Processing
We used the Re-DocRED dataset for fine-tuning, refining it by removing duplicates and ensuring factual accuracy. This involved adjusting reciprocal relations like “follows” and “followed by” to accurately represent inversion, enhancing the dataset’s robustness and precision.
@@
Addressing this, we systematically rewrote all 96 relation descriptions, markedly improving model performance in Table 2. An example of our revised description is as follows. Details of all relation descriptions are presented in Table 6.
@@
Finally, in line with the RHF paradigm, we crafted instruction fine-tuning templates, breaking down the RE process of each sample into three distinct steps. The specific details of these templates can be found in Table 7, we provide a display of how we constructed our training data using simple prompt templates, the extraction of relations, the extraction of the head entity, and finally the triplet extraction.
@@ 4.1 Experimental Setup
Model dev F1 test F1
TAG 49.34 49.38
AutoRE-ChatGLM3-6B 49.86 51.11
AutoRE-Mistral-7B 53.01 51.91
AutoRE-Vicuna-7B 54.29 53.84
Table 4: The results of AutoRE for different PLMs. Compared with TAG, all AutoRE models achieve the best performance.
@@
[…]
[width=.85]figures/Ablation.pdf
4.2 Overall Result
After training three distinct QLoRA modules, we test the performance on the Re-DocRED and then combine three QLoRAs to get the final performance on the dev set and test set, the result is shown in Table 3. When compared with TAG (Zhang et al., 2023) as a baseline which firstly reported the end-to-end RE on Re-DocRED, our method has achieved SOTA results, as shown in Table 4. In both the dev set and test set, the performance improvement of AutoRE finetuned with Mistral-7B over TAG is approximately 7.44% on the dev set, and about 5.12% on the test set demonstrating the effectiveness of our approach. Furthermore, by decomposing the task into three subtasks and training with three LoRA modules, we not only achieved excellent results but also naturally acquired an easily extendable trait. This allows for targeted improvement of a specific module’s performance without impacting the performance of other subtasks. Additionally, it is worth noting that our work is the first to utilize large language models for processing the Re-DocRED dataset. AutoRE can serve as a reference for subsequent research in this field.
@@ 4.3 Ablation
In the ablation study, we employed Mistral-7B to fine-tune the paradigms mentioned before, revealing that the RHF model yields the best performance when solely utilizing one QLoRA module. This finding substantiates our initial hypothesis during paradigm selection: employing a step-by-step approach enhances the extraction of triplet facts while significantly reducing erroneous triplets through fine-tuning. Building on this, we compared the impact of including descriptions versus omitting them. The results confirmed that incorporating proper relation descriptions indeed benefits the model, as shown in Figure 4. Additionally, we explored the effectiveness of training the entire dataset with one QLoRA versus independently training different stages of RHF with three distinct QLoRAs. The latter approach demonstrated superior performance. We believe this is due to the data imbalance among predicting relations, predicting head entities, and predicting factual triples in the RHF paradigm, with the data volume for the three subtasks begin approximately 2.8%, 24.23%, and 72.97%, respectively. When combined, the model tends to favor the prediction of triples, while its capability to predict relations is relatively insufficient.
Additionally, we have applied this framework to Vicuna-7B and ChatGLM3-6B, and both models surpassed the current SOTA levels, demonstrating the universality of the AutoRE framework. The comparative results of these experiments are illustrated in the accompanying Figure 4. Vicuna-7B scored the highest, surpassing TAG by 10.03% and 9.03% respectively on the dev and test set, whereas ChatGLM3-6B was somewhat lower. This might be due to ChatGLM3-6B having a higher proportion of Chinese in its pre-training, while it was tested on an English task. We have deployed the system on the online platform [^5] https://models.aminer.cn/neptune/ for users to access and experience, as shown in Figure 3.
@@ 5 Conclusion
Acknowledgements
This work is supported by Technology and Innovation Major Project of the Ministry of Science and Technology of China under Grant 2022ZD0118600, Natural Science Foundation of China(NSFC) 62276148 and 62425601, the New Cornerstone Science Foundation through the XPLORER PRIZE.